# Cosmic Ray Diffusion in the Turbulent Interstellar Medium: Effects of Mirror Diffusion and Pitch Angle Scattering

### v1

@@ 4 Discussion
Figure 10: The same as in Figure 9, but considering a smoothed background distribution of the magnetic field (as in Figure 8).
In this work we have explored the effects of mirror diffusion of CRs by combining 3D MHD simulations of a supersonic, sub-Alfvenic turbulent molecular cloud a with test particle simulations, considering initial pitch angles corresponding to $\mu=0.2$ and $0.8$.
Particles injected into the turbulent domain undergo both bouncing and scattering processes. We have observed that mirror diffusion significantly influences their parallel diffusion component relative to the mean magnetic field, validating the LX21 predictions.
Analyzing their $\mu_{b}$, defined as the last maximum value of $\mu$ where their adiabatic invariant $\zeta$ exhibits minimal variation over time, reveals that the majority of bounces occur for $\mu_{b}<0.6$, however this should depend on the background magnetic field.
The initial diffusion of CRs is influenced by their initial pitch angle, with those starting closer to $90^{\circ}$ (or $\mu_{0}$ closer to 0) being more prone to mirror diffusion and requiring longer diffusion times, while those with smaller pitch angles (larger $\mu_{0}$) can escape the system more rapidly. This disparity also affects the time required for particles to attain normal diffusion. Particles beginning with pitch angles near $90^{\circ}$ achieve a nearly constant $\Delta\tilde{z}^{2}/\Delta t$ at a slower rate (see Figures 7 and 8). Suppression of pitch angle scattering can further accentuate the role of initial pitch angle in defining the particle diffusion regime, depending on its Larmor radius.
@@
We can convert the parallel diffusion coefficient $D_{\parallel}$ (the value of $\Delta\tilde{z}^{2}/\Delta t$ after it achieves a nearly stable value), obtained from the simulations in code units, into physical units, considering an arbitrary physical size for the system, as follows:
$\begin{split}D_{\parallel,phys}[cm^{2}/s]&=D_{\parallel}\cdot L_{0}\,c\\ &=D_{\parallel}\cdot 2.776\times 10^{29}\,\Big{(}\frac{L_{0}}{3pc}\Big{)},\end{split}$ (7)
@@
$\begin{split}B[\mu G]=3.215\,\Big{(}\frac{n}{10cm^{-3}}\Big{)}^{\frac{1}{2}}\Big{(}\frac{c_{s}}{1.9\cdot 10^{4}cm/s}\Big{)}\\ \Big{(}\frac{M_{s}}{7.0}\Big{)}\Big{(}\frac{M_{A}}{0.6}\Big{)}^{-1},\end{split}$ (8)
@@
$R_{L}[pc]$ $=$ $1.084\frac{(E/1PeV)}{(B/1\mu G)}\Big{(}\frac{v_{\perp}}{c}\Big{)},$ (9)
$\Omega[s^{-1}]$$\approx$ $9.743\times 10^{-9}\Big{(}\frac{1pc}{R_{L}}\Big{)},$ (10)
where $E$ is the energy of the particle, $\Omega$ its gyrofrequence, and $R_{L}$ its Larmor radius in second and pc units, respectively. As stressed before, we have assumed $v_{\perp}\approx c$.
As emphasized in Section 1, recent findings from VHE gamma-ray observatories like H.E.S.S., HAWC and LHAASO have revealed extensive diffuse emissions reaching a few to hundreds of TeV and even PeV within our galaxy (Abeysekara et al., 2017; Huang et al., 2018; Abeysekara et al., 2021; Cao et al., 2021), without definitive identification of their sources. Consequently, comprehending CR acceleration and diffusion in the surrounds of embedded sources has become critically important for interpreting the observations. Traditional models and assumptions to characterize CR diffusion are insufficient to elucidate these emissions.
@@
We note that $D_{\parallel,phys}$ is consistent with the implied values to explain VHE observations of extended sources like Geminga and PSR B0656+14 (Abeysekara et al., 2017). Their combined halo has an inferred diffusion coefficient of $4.5\times 10^{27}cm^{2}/s$ for 100 TeV CRs within an extended zone with tens of parsec (Abeysekara et al., 2017). Similarly, Huang et al. (2018) propose a diffusion coefficient $\lesssim 10^{28}\,cm^{2}/s$ for electrons of $\sim 10$TeV in a region of a few tens of parsecs around Vela X.
@@
Interestingly, our results from the models featuring smoothed background magnetic field distribution (Figures 8 and 10) could also offer insights into systems like the nuclear regions of starburst galaxies. For example, Krumholz et al. (2020) argues that, as in molecular clouds within our Galaxy, these systems are likely to possess sufficiently low ionization fractions. Consequently, ion-neutral damping would halt turbulent cascades at scales much larger than the gyroradius of CRs with energies in the range of a few hundred TeV. Still, CRs could scatter off turbulence self-generated by streaming instability. The competition between the streaming instability and ion-neutral damping leads to a CR streaming speed equal to the ion Alfvén speed, which is independent of CR energy up to $\sim 1TeV$. Diffusion should occur only due to field line random walk along the magnetic field, yielding a diffusion coefficient of approximately $10^{27}-10^{28}\,\text{cm}^{2}/\text{s}$, in this energy range. While the behavior of these CRs would align with the results shown in Figures 8 and 10 in the absence of streaming instability feedback, proper modeling of diffusion in such environments necessitates the incorporation of CR feedback effects too.
Considering all the preceding discussions, it’s worth noting that while pitch angle scattering alone fails to generate sufficiently small diffusion coefficients to align with observations, mirror diffusion alone also falls short. However, the interplay between both mechanisms appears to yield the slowest diffusion coefficients. This is evident when examining, for instance, all models with $R_{L}/L_{0}=0.03$ (indicated by blue curves) in Figures 9 and 10. In the former, $D_{\parallel}$ stabilizes around $\sim 10^{27}\,\text{cm}^{2}/\text{s}$, whereas in the latter, it continues to increase, without yet displaying indications of approaching a nearly constant value by the end of the considered timeframe.
The variation of $\mu$ due to resonance scattering allows for more particles to interact with mirrors, resulting in a net coefficient that is smaller than the one obtained from each process separately. Therefore, mirror diffusion needs to be taken into account with pitch angle scattering to properly model CR diffusion.
@@
Lastly, it is worth noting that recent studies by Lemoine (2023) and Kempski et al. (2023) have examined the implications of intermittent bends and reversals in magnetic fields on CR diffusion. They conclude that low-energy particles experience more efficient confinement than high-energy particles, attributed to the absence of pitch angle scattering and isotropization. While employing different configurations for the background magnetic field — such as utilizing incompressible turbulence in the former or omitting a guide field in the latter — their findings exhibit similarities to the outcomes observed in our models, especially in the ones with smoothed background distribution, particularly regarding $R_{L}/L_{0}=0.03$. The absence of pitch angle scattering is a significant factor influencing fast particle escape rates ($\mu_{0}=0.8$) or prolonged mirroring ($\mu_{0}=0.2$). However, further investigation is warranted to explore the intersections between their studies and ours.
@@ 5 Conclusions
In this work, we have numerically demonstrated the process of mirror diffusion theoretically predicted in LX21. This effect is particularly significant for particles with large pitch angles. We found that in combination with resonance scattering, the particle transport is akin to a Levy-flight-like propagation. Our numerical results agree with those in Zhang & Xu (2023), but extended them accounting for both mirror and scattering diffusion to assess their combined effects aiming at applications to realistic astrophysical conditions.
We have found that when pitch angle scattering is minimized, the impact of mirror diffusion depends on the particle’s initial pitch angle, for a significant amount of time. This effect is especially evident for lower energy particles.
Furthermore, our findings have revealed that the range of pitch angles subjective to mirroring aligns with the predictions derived from the models of Xu & Lazarian (2020) and LX21, and the estimated physical values of the diffusion coefficient, $\lesssim 10^{27}cm^{2}/s$, fall within recent constraints imposed by VHE energy observations.
The combination of resonance scattering with mirror diffusion leads to a slower diffusion compared with the case with scattering alone of particles and we contend that mirror diffusion should be considered when studying CR propagation within turbulent magnetized environments. Moving forward, we intend to integrate these effects into comprehensive simulations of CR propagation in massive star-forming regions in order to quantify their VHE emission.
L.B-M. acknowledges support from the Brazilian funding agency CAPES, CNPq and also the kind hospitality of Professor Alex Lazarian and the University of Wisconsin - Madison; E.M.G.D.P. acknowledges support from the Brazilian funding agencies FAPESP (grants 2013/10559-5 and 2021/02120-0) and CNPq (grant 308643/2017-8); This research was also supported in part by grant no. NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). The simulations presented in this work were performed in the cluster of the Group of Plasmas and High-Energy Astrophysics of IAG-USP (GAPAE) acquired with support from FAPESP (grant 2013/10559-5).

### v3

@@ 4 Discussion
Figure 10: The same as in Figure 9, but considering a smoothed background distribution of the magnetic field (as in Figure 8), and a much larger time interval. Models with $R_{L}=0.03L_{0}$ take a longer time to saturate due to the reduced scattering. See text for further details.
Figure 11: Distribution of $\mu=\cos{\theta}$ at $t\cdot\Omega=10^{4}$ (or $t=4.6\cdot 10^{13}s$) for the smoothed models with $R_{L}/L_{0}=0.03$ (blue lines in Figure 10).
In this work we have explored the effects of mirror diffusion of CRs by combining 3D MHD simulations of a supersonic, sub-Alfvenic turbulent molecular cloud with test particle simulations, considering initial pitch angles corresponding to $\mu=0.2$ and $0.8$. We considered two background magnetic field distributions - the smoothed and the complete models - for CR propagation analysis. The study of the smoothed model was entailed only to highlight the effects of mirror diffusion over scattering. Nevertheless, in realistic astrophysical media, the damping scale of MHD turbulence can be larger or smaller than the gyroradius of the CRs under consideration. Therefore, both models investigated (smoothed and complete) are realistic.
Particles injected into the turbulent domain undergo both bouncing (or mirroring) and resonant scattering processes. We have observed that mirror diffusion significantly influences their parallel diffusion component relative to the mean magnetic field, which is consistent with LX21 predictions. Turbulent magnetic mirrors can still happen.
Analyzing their $\mu_{b}$, defined as the last maximum absolute value of $\mu$ where their adiabatic invariant $\zeta$ exhibits minimal variation over time, reveals that the majority of bounces occur for $\mu_{b}<0.6$, however, this should depend on the background magnetic field.
The initial diffusion of CRs is influenced by their initial pitch angle, with those starting closer to $90^{\circ}$ (or $\mu_{0}$ closer to 0) being more prone to mirror diffusion and requiring longer diffusion times, while those with smaller pitch angles (larger $\mu_{0}$) can escape the system more rapidly. This disparity also affects the time required for particles to attain normal diffusion. Particles beginning with pitch angles near $90^{\circ}$ achieve a nearly constant $\Delta\tilde{z}^{2}/\Delta t$ at a slower rate (see Figures 7 and 8). Suppression of pitch angle scattering can further accentuate the role of the initial pitch angle in defining the particle diffusion regime, depending on its Larmor radius.
@@
$\begin{split}D_{\parallel,phys}[\mathrm{cm^{2}/s}]&=D_{\parallel}\cdot L_{0}\,c\\ &=D_{\parallel}\cdot 2.776\times 10^{29}\,\Big{(}\frac{L_{0}}{3\,\mathrm{pc}}\Big{)},\end{split}$ (7)
@@
$\begin{split}B[\mathrm{\mu G}]=3.215\,\Big{(}\frac{n}{10\,\mathrm{cm^{-3}}}\Big{)}^{\frac{1}{2}}\Big{(}\frac{c_{s}}{1.9\cdot 10^{4}\,\mathrm{cm/s}}\Big{)}\\ \Big{(}\frac{M_{s}}{7.0}\Big{)}\Big{(}\frac{M_{A}}{0.6}\Big{)}^{-1},\end{split}$ (8)
@@
$R_{L}[\mathrm{pc}]$ $=$ $1.084\frac{(E/1\mathrm{PeV})}{(B/1\mathrm{\mu G})}\Big{(}\frac{v_{\perp}}{c}\Big{)},$ (9)
$\Omega[\mathrm{s^{-1}}]$$\approx$ $9.743\times 10^{-9}\Big{(}\frac{1\mathrm{pc}}{R_{L}}\Big{)},$ (10)
where $E$ is the energy of the particle, $\Omega$ its gyrofrequency, and $R_{L}$ its Larmor radius in second and pc units, respectively. As stressed before, we have assumed $v_{\perp}\approx c$.
As emphasized in Section 1, recent findings from VHE gamma-ray observatories like H.E.S.S., HAWC and LHAASO have revealed extensive diffuse emissions reaching a few to hundreds of TeV and even PeV within our galaxy (Abeysekara et al., 2017; Huang et al., 2018; Abeysekara et al., 2021; Cao et al., 2021), without definitive identification of their sources. Consequently, comprehending CR acceleration and diffusion in the surroundings of embedded sources has become critically important for interpreting the observations. Traditional models and assumptions to characterize CR diffusion are insufficient to elucidate these emissions.
@@
We note that $D_{\parallel,phys}$ is consistent with the implied values to explain VHE observations of extended sources like Geminga and PSR B0656+14 (Abeysekara et al., 2017). Their combined halo has an inferred diffusion coefficient of $4.5\times 10^{27}\mathrm{cm^{2}/s}$ for 100 TeV CRs within an extended zone with tens of parsec (Abeysekara et al., 2017). Similarly, Huang et al. (2018) propose a diffusion coefficient $\lesssim 10^{28}\,\mathrm{cm^{2}/s}$ for electrons of $\sim 10$ TeV in a region of a few tens of parsecs around Vela X. Fang (2022) provides a comprehensive review of the characteristics of recent PWNe VHE observations and possible scenarios to explain the small diffusion coefficients which are also compatible with our findings. We refer to his work for more detailed information.
@@
Interestingly, our results from the models featuring smoothed background magnetic field distribution (Figures 8 and 10) could also offer insights into systems like the nuclear regions of starburst galaxies. For example, Krumholz et al. (2020) argues that, as in molecular clouds within our Galaxy, these systems are likely to possess sufficiently low ionization fractions. Consequently, ion-neutral damping would halt turbulent cascades at scales much larger than the gyroradius of CRs with energies in the range of a few hundred TeV. Still, CRs could scatter off turbulence self-generated by streaming instability. The competition between the streaming instability and ion-neutral damping leads to a CR streaming speed equal to the ion Alfvén speed, which is independent of CR energy up to $\sim 1\mathrm{TeV}$. Diffusion should occur only due to field line random walk along the magnetic field, yielding a diffusion coefficient of approximately $10^{27}-10^{28}\,\mathrm{cm^{2}/s}$, in this energy range. While the behavior of these CRs would align with the results shown in Figures 8 and 10 in the absence of streaming instability feedback, proper modeling of diffusion in such environments necessitates the incorporation of CR feedback effects too.
Resonant scattering occurs only if the resonance condition is satisfied. For the resonant wavenumber $k_{res}$, we can write:
$k_{res}2\pi R_{L}\simeq\frac{(1-\cos^{2}{\theta_{res}})^{0.5}}{\cos{\theta_{res}}},$ (11)
where $\theta_{res}$ is the angle between the particle’s velocity and the magnetic field that satisfies the resonance condition. In our smoothed models, smoothing limits $k_{res}$ to values below 5. For $R_{L}=0.03L_{0}$, we find $\tan{\theta_{res}}\sim 0.9425$ or $\mu_{res}\sim 0.73$ ($\theta_{res}\sim 43^{\circ}$). This implies that resonant scattering will be dominant only for approximately $\mu\gtrsim 0.73$. Referring to Figure 5, all bounces occur with $\mu_{c,\text{max}}<0.7$ (panels (c) and (d)). However, some particles eventually cease bouncing and align more closely with the magnetic field, reaching $\mu\gtrsim 0.73$. This transition triggers resonant scattering, enabling particles to diffuse more quickly across the system (see Figure 4, see also Appendix A). This mechanism explains the higher diffusion coefficient observed in the smoothed case compared to the complete magnetic field distribution for $R_{L}=0.03L_{0}$ (Figures 7 and 9).
Indeed, in the complete case, mirroring remains dominant for $\mu\lesssim 0.7$, but scattering occurs across all $\mu$ values. Consequently, in a fully developed magnetic field, the interplay between these mechanisms at large pitch angles becomes more frequent starting already at small spatial scales, leading to more transitions. These transitions slow diffusion significantly compared to the smoothed scenario.
This behavior is evident in all models with $R_{L}/L_{0}=0.03$ (indicated by blue curves). Comparing diffusion coefficients in Figures 9 and 10, we observe that in the complete magnetic field distribution, $D_{\parallel}$ stabilizes around $\sim 10^{27}\text{cm}^{2}/\text{s}$, whereas in the smoothed case, $D_{\parallel}$ continues to increase by nearly an order of magnitude more, with a distinct trend of delayed saturation.
For $R_{L}=0.06L_{0}$, the resonant condition gives $\theta_{res}\lesssim 63^{\circ}$ or $\mu_{res}\gtrsim 0.46$. For $R_{L}=0.10L_{0}$, $\theta_{res}\lesssim 73^{\circ}$ or $\mu_{res}\gtrsim 0.30$. This means that for $R_{L}=0.03L_{0}$, the range of $\mu$ values subject to scattering is narrower. Particles with larger $R_{L}$ transition more readily between the two regimes, explaining their behavior, which resembles that of the complete magnetic field distribution models. While both scattering and mirroring are present in all cases, $R_{L}=0.03L_{0}$ provides the clearest distinction between these processes.
The discussion above can be further stressed. It is important to note that suppressing small-scale structures in the smoothed models (Figures 8 and 10) not only reduces the effectiveness of pitch-angle scattering but also eliminates small-scale magnetic mirrors, as discussed in Section 3.2. These small-scale mirrors are more effective at reflecting particles due to their stronger longitudinal magnetic field gradients compared to larger-scale mirrors. In Figures 8 and 10, the scattering is reduced for the blue curve. At early times, diffusion with $\mu_{0}=0.2$ is slower than that with $\mu_{0}=0.8$ due to the stronger mirroring at small $\mu$. Later in Figure 10, as some particles experience an increase in $\mu$—despite inefficient scattering—the reduced scattering at large $\mu$ accelerates overall diffusion, resulting in faster transport. Thus, in the smoothed magnetic field, diffusion becomes more efficient compared to the complete field case, reinforcing the conclusions presented earlier.
The mean free path in scattering regime alone at small pitch angles would be infinity. However, in reality, it is not possible to study scattering regime in isolation, as scattering inevitably causes particles to transit between scattering and mirroring regimes - i.e., small to large pitch angles (and vice versa). Our study with a smoothed magnetic field (which still includes weak scattering) clearly shows this trend (see the blue line in Figure 10). In the case of mirroring, it is possible to separately study its effect on diffusion near the source of CRs, as scattering (small pitch-angle) CRs escape rapidly, leaving behind only mirroring (large pitch-angle) CRs near the source. This entails an enhancement of CRs, due to their slow mirror diffusion, and gamma rays specially near the source, an effect that is observable (e.g. Abeysekara et al., 2017).
@@
It is worth noting that recent studies by Lemoine (2023) and Kempski et al. (2023) have examined the implications of intermittent bends and reversals in magnetic fields on CR diffusion. They conclude that low-energy particles experience more efficient confinement than high-energy particles, attributed to the absence of pitch angle scattering and isotropization. All their models consider very small Larmor radius, and the only model among ours that has a similar initial configuration for the particles is the one with a smoothed background magnetic field with $R_{L}/L_{0}=0.03$. However, while we consider supersonic and sub-Alfvénic turbulence, those works employ different background magnetic field configurations. The reduced influence of pitch angle scattering is a significant factor influencing fast particle escape rates ($\mu_{0}=0.8$) or prolonged mirroring ($\mu_{0}=0.2$). However, further investigation is warranted to explore the intersections between their studies and ours. More recently, Zhang & Xu (2024) have investigated the influence of mirror diffusion in a similar scenario to that employed by Lemoine (2023) and Kempski et al. (2023). We refer to their work for a comparison between these different processes.
Lastly, a recent study by Bell et al. (2025) demonstrated that magnetic mirroring, combined with modest small-angle scattering, can reduce cosmic ray transport to levels approaching Bohm diffusion. Their Vlasov-Fokker-Planck analysis shows how large-scale magnetic mirrors effectively confine CRs near shocks, potentially enabling acceleration to PeV energies without the need for magnetic field amplification on the Larmor scale. By bridging the gap between idealized diffusion models and realistic magnetic field geometries, their work—complementary to the present study—reinforces the need to incorporate mirroring effects in theories of CR acceleration and propagation.
@@ 5 Conclusions
In this work, we have numerically demonstrated the process of mirror diffusion theoretically predicted in LX21. This effect is particularly significant for particles with large pitch angles. We found that in combination with resonance scattering, the particle transport is akin to a Levy-flight-like propagation. Our numerical results agree with those in Zhang & Xu (2023), but extended them accounting for both mirror and resonant scattering diffusion to assess their combined effects aiming at applications to realistic astrophysical conditions.
Mirror effects are essential for small $\mu$ propagation (large pitch angle), and so they are relevant while addressing the $\sim 90-$degree scattering problem and its impact on the effective diffusion.
We have found that when pitch angle resonant scattering is significantly suppressed, the impact of mirror diffusion depends on the particle’s initial pitch angle, for a significant amount of time. This effect is especially evident for lower energy particles.
Furthermore, our findings have revealed that the range of pitch angles subjective to mirroring aligns with the predictions derived from the models of Xu & Lazarian (2020) and LX21, and the estimated physical values of the diffusion coefficient, $\lesssim 10^{27}\mathrm{cm^{2}/s}$, fall within recent constraints imposed by VHE energy observations.
Resonant scattering at small pitch angles alone would lead to an infinite particle mean free path, as it cannot reverse the direction of particle motion. The combination of resonance scattering with mirror diffusion leads to a slower diffusion compared with predictions that consider scattering alone of particles (see for a discussion e.g. Felice & Kulsrud, 2001; Shalchi et al., 2009; Delahaye et al., 2010; Cohet & Marcowith, 2016; Abeysekara et al., 2017; Krumholz et al., 2020, and references therein), and we contend that mirror diffusion should be considered when studying CR propagation within turbulent magnetized environments. Moving forward, we intend to integrate these effects into comprehensive simulations of CR propagation in massive star-forming regions in order to quantify their VHE emission.
L.B-M. acknowledges support from the Brazilian funding agency CAPES, CNPq and also the kind hospitality of Professor Alex Lazarian and the University of Wisconsin - Madison; E.M.G.D.P. acknowledges support from the Brazilian funding agencies FAPESP (grants 2013/10559-5 and 2021/02120-0) and CNPq (grant 308643/2017-8). SX acknowledges the support from NASA ATP award 80NSSC24K0896. A.L. acknowledges the support of 823 wNSF grants AST 2307840. This research was also supported in part by grant no. NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP), and by the Aspen Center for Physics through the NSF grant PHY-2210452 and Durand Fund. The simulations presented in this work were performed in the cluster of the Group of Plasmas and High-Energy Astrophysics of IAG-USP (GAPAE) acquired with support from FAPESP (grant 2013/10559-5). We also acknowledge inspiring and very fruitful discussions with Anthony Bell, Philipp Kempski, Kun Fang and the participants of the program on “Turbulence in Astrophysical Environments” at KITP.
@@
Appendix A Pitch angle scattering and large particle displacement
Figure 12: Particle’s pitch-angle evolution and displacment parallel to the magnetic field along time. Initially the particle has $\mu_{0}=0.2$, $R_{L}/L_{0}=0.03$, and a smoothed background magnetic field distribution. Orange regions around the blue line show parts of the trajectory where the estimated condition for scattering is satisfied.
Figure 12 illustrates the trajectory of a particle characterized by $\mu_{0}=0.2$, $R_{L}=0.03L_{0}$, within a smoothed background magnetic field. The orange circles highlight segments of the trajectory where $\mu>0.73$, which, as discussed in Section 4, corresponds to the scattering regime for particles in this type of magnetic field background. This example demonstrates that the particle exhibits the most significant diffusion across the system when it is in the scattering regime.
Appendix B Alternative method to identify bounces
Figure 13: Distribution of $\zeta_{c}/\zeta_{avg}$ for all bounce candidates in each simulation. The black dashed line shows a Gaussian distribution fit to the characteristic peak around $\zeta_{c}/\zeta_{avg}=1$ for models with $R_{L}=0.03L_{0}$.
Figure 14: Same as Figures 5 and 6 but using Gaussian method for the identification of $\mu_{b}$.
Table 4: Same as Table 2 but using the Gaussian method to identify mirror bounces.
Complete $\mu_{0}=0.2$ $\mu_{0}=0.8$
$R_{L}$ Mirror Bounce Total % Mirror Bounce Total %
0.03 91361 123859 73.76 89931 120331 74.74
0.06 165459 223060 74.18 161250 218128 73.92
0.10 246912 322056 76.67 234385 307357 76.26
Smooth $\mu_{0}=0.2$ $\mu_{0}=0.8$
$R_{L}$ Mirror Bounce Total % Mirror Bounce Total %
0.03 247899 354073 70.01 110200 152117 72.44
0.06 126401 154984 81.56 114512 136214 84.07
0.10 108227 140372 77.10 105874 137993 76.72
We propose here an alternative method to identify bounces to the one introduced in Section 3.1. We first analyze the distribution of $\zeta_{c}/\zeta_{avg}$ for all bounce candidates. As illustrated in Figure 13, a peak emerges near $\zeta_{c}/\zeta_{avg}=1$, coinciding with the $\zeta/\zeta_{avg}$ values where the majority of bounces occur. We then fit a Gaussian distribution to the peak and classify all values within $1.8\sigma_{fit}$ of the peak center as mirror bounces.
The results obtained using this method are largely consistent with those presented in Section 3.2. However, some differences also arise with regard to Figures 5 and 6. With this new approach, most of the data points below the orange line in Figure 5 are eliminated, while in Figure 6, the tail of each distribution decays to zero for $\mu_{b}>0.8$. See Figure 14 for comparison.
Although this method seems to improve the results, there is still a selection effect. We cannot guarantee that changes in direction which are not associated to the peak around $\zeta_{c}/\zeta_{avg}=1$, are not due to mirrors. In other words, we could have mirror events with smaller values of $\zeta_{c}/\zeta_{avg}$.
Panel (d) of Figure 13 highlights another important feature: a distinct secondary peak near $\zeta_{c}/\zeta_{avg}=0$ for $R_{L}=0.03L_{0}$. While this might initially suggest an increased role of scattering—since $\zeta$ approaches zero as $\mu\rightarrow 1$—it is important to consider the nature of particle motion. In the absence of significant scattering, particles move ballistically along magnetic field lines with conserved magnetic moment, a process known as wandering diffusion (e.g., Zhang & Xu, 2024). To differentiate between wandering diffusion and mirroring diffusion: the former occurs at large, nearly constant $\mu$ values (indicating no change in direction relative to the local magnetic field), whereas the latter involves small, oscillating $\mu$ with frequent crossings through $\mu=0$. A closer analysis shows that the Gaussian distribution accounts for over 72% of all candidate bounces, pointing to mirroring as the dominant mechanism. Even though the selected conditions in this case favor wandering diffusion, mirroring still emerges as the primary cause of particle deceleration. In all other cases examined, mirror-induced bounces are even more clearly dominant.
Table 4 compares the total number of identified events. While the percentage of mirror events improves relative to Table 2 in most cases, the true value likely lies between the two estimates, considering the discussion above. Also, the slightly smaller percentages for smaller Larmor radius rather than larger ones, are directly related to the secondary peaks in Figure 13. Nevertheless, the analysis consistently supports the conclusion that mirroring is the dominant mechanism in CR diffusion.